Find answers to the following. Write and indicate how you solved them.
 Is 5 / 9 equal to 4 / 5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Convert these fractions into like fractions
5 / 9 = (5 / 9)× (5 / 5)
= 25 / 45
4 / 5 = (4 / 5) × (9 / 9)
= 36 / 45
∴25 / 45 ≠ 36 / 45
Hence, 5 / 9 is not equal to 4 / 5

​Thanks